The length of a rectangle is five times its width.
If the perimeter of the rectangle is 64m , what is the area

Answer by mananth(16946)

let width be x
length = 5x
..
perimeter P = 2L+2W
P= 2(x+5x)
P=12x
12x=64
x= 5.33 feet width
length = 5*5.33= 26.65
..
Area = 26.65 * 5.33
= 142 sq.m